OpenStudy (anonymous):

A region R, in the first quadrant only, is enclosed by y=x^3 and y=cubed root of x. find the volume of the solid obtained by revolving the region R about x=-2.

OpenStudy (anonymous):

First you have to imagine and draw a rough of what these would look like. You would find examples of these functions in the first chap of a calculus book or put them in a calculator

OpenStudy (anonymous):

okay i imagined it and drew a picture. what do i do next?

OpenStudy (anonymous):

Then you find the boundaries of integration (it might be obvious, it starts at 0) but set x^3=cube root of x; solve for x. The points tell you where y=x^3 and y=cube root x intersect (boundaries of integration)

OpenStudy (anonymous):

ehhh i tried but i dont know how to find the boundaries. i got x^3-sq root of x=0 but i dont know what that would make x equal?

OpenStudy (anonymous):

x^3=cube root of x The problem is the cube root sign. Just cube each side of the eq (x^3)^3=(cube root of x)^3

OpenStudy (anonymous):

okay i got x^9-x^2

OpenStudy (anonymous):

how do i factor that?

OpenStudy (anonymous):

Good. but you are off a little. It should be x^9-x=0 Both x^9 and x have a common factor of x, so you have to pull it out

OpenStudy (anonymous):

but if i square the squareroot of x, isnt is x^2?

OpenStudy (anonymous):

Actually if you square the square root of x, you get x; likewise cube the cube root you get x

OpenStudy (anonymous):

ohhhh okay sorry i got it!

OpenStudy (anonymous):

So, you have a lot of catching up, I'm going to go fast because I have to go. So the boundaries are 0 to 1

OpenStudy (anonymous):

okay next?

OpenStudy (anonymous):

You are going to be using washer method

OpenStudy (anonymous):

Radius of big washer is x^3 +2 radius of small washer is cube root of x +2

OpenStudy (anonymous):

okay ill try that!

OpenStudy (anonymous):

Integrate from 0 to 1 [{(x^3+2)^2}pi - {(cube rooot x +2)^2}pi]

OpenStudy (anonymous):

okay i will do that now. whats the final answer?

OpenStudy (anonymous):

Haven't work it out. Setting it up is the fun part. Integrating is the dirty work. I left the dirty work for you.

OpenStudy (anonymous):

okay thanksss for all the help!